# Facility Location Games with Scaling Effects

### v1

@@ 6.2 Maximum Cost
We next prove approximation ratio results for the objective of maximum cost, a standard measure of egalitarian fairness. We begin by giving the maximum cost approximation ratio of any phantom mechanism with $p_{1}=0$ and $p_{n+1}=1$, and again the proof applies to both continuous and piecewise linear scaling functions.
@@
For continuous and piecewise linear scaling functions, any phantom mechanism with $p_{1}=0$ and $p_{n+1}=1$ has a maximum cost approximation ratio of $2r_{q}$.
@@
We first prove that for every phantom mechanism $f$ with $p_{1}=0$ and $p_{n+1}=1$, there exists an agent profile where the facility is placed at either $x_{1}$ or $x_{n}$. Due to symmetry, it suffices to consider the former case, where we require that $p_{2}<1$ and $x_{1}=x_{2}=\dots=x_{n-1}>p_{2}$. Consider such an agent location profile where the facility is placed at $x_{1}$ and $x_{1}\neq x_{n}$, and also consider a scaling function which has a global maximum at $x_{1}$ and a global minimum at $\frac{x_{1}+x_{n}}{2}$. We have $MC(q,f(\boldsymbol{x}),\boldsymbol{x})=(x_{n}-x_{1})\max_{y\in X}q(y)$ and $MC(q,y^{*}_{MC},\boldsymbol{x})=\frac{x_{n}-x_{1}}{2}\min_{y\in X}q(y)$, giving $\frac{MC(q,f(\boldsymbol{x}),\boldsymbol{x})}{MC(q,y^{*}_{MC},\boldsymbol{x})}=2r_{q}$. Thus, for every phantom mechanism with at least one phantom on 00 and $1$, we can construct an agent location profile and scaling function such that the cost ratio is $2r_{q}$, proving the lower bound.
To see that the approximation ratio is at most $2r_{q}$, note that a phantom mechanism with $p_{1}=0$ and $p_{n+1}=1$ ensures the facility is placed within $[x_{1},x_{n}]$, so the mechanism’s maximum cost is at most $(x_{n}-x_{1})\max_{y\in X}q(y)$, whilst the optimal maximum cost is at least $\frac{x_{n}-x_{1}}{2}\min_{y\in X}q(y)$.
@@
Consider an agent location profile where $x_{1}=x_{2}=\dots=x_{n-1}$ and $x_{n}\neq x_{1}$. When the scaling function has its minimum value at $\frac{x_{1}+x_{n}}{2}$ and its maximum value on the facility/dictator, we have $\frac{MC(q,f(\boldsymbol{x}),\boldsymbol{x})}{MC(q,y^{*}_{MC},\boldsymbol{x})}=2r_{q}$, showing that the approximation ratio is at least $2r_{q}$. For the matching upper bound, we note that the mechanism’s maximum cost is at most $(x_{n}-x_{1})\max_{y\in X}q(y)$ and the optimal maximum cost is at least $\frac{x_{n}-x_{1}}{2}\min_{y\in X}q(y)$.

### v2

@@ 7.2 Maximum Cost
We next prove approximation ratio results for the objective of maximum cost, a standard measure of egalitarian fairness. We begin by giving the maximum cost approximation ratio of any phantom mechanism with scaling with $p_{1}=0$ and $p_{n+1}=1$, and again the proof applies to any arbitrary continuous scaling function.
@@
Any phantom mechanism with scaling has a maximum cost approximation ratio of $2r_{q}$.
@@
Similar to the proof of Theorem 7.3, given a scaling function $q$, we consider two cases: either all phantoms are located at 00 or $1$, or there is at least one phantom that is not located at 00 or $1$. In the former case, where the facility is located at some agent, we can always construct an agent location profile where the scaling function is highest at the agent that receives the facility, so the corresponding maximum cost must be at least $\max_{y}q(y)\sum_{i}|f_{mid}(\boldsymbol{x})-x_{i}|$, where $f_{mid}$ is the midpoint mechanism. Now suppose that the location profile is also constructed such that the scaling function is lowest at $(x_{1}+x_{n})/2$. Placing the facility at this point results in the optimal maximum cost of $\min_{y}q(y)\sum_{i}|f_{mid}(\boldsymbol{x})-x_{i}|$, and thus the total cost approximation ratio must be at least $2r_{q}$ in this case.
For the case where at least one phantom is not located at 00 or $1$, by Lemma 2, it suffices to consider the phantom mechanism with scaling $f^{\prime}$, which is defined in Lemma 2 and the proof of Theorem 7.3. We show that the approximation ratio of $f^{\prime}$ is at least $2r_{q}$. Consider a scaling function with two global minima at $c_{1}$ and $c_{3}$, and a global maximum at $c_{2}$, where $0<c_{1}<c_{2}<c_{3}$. The $n-1$ phantoms of $f^{\prime}$ are placed at $c_{1}$. We construct an agent location profile such that $x_{1}=\dots=x_{n-1}=c_{2}$, and $x_{n}=2c_{3}-c_{2}$. Here, $f^{\prime}$ places the facility at $c_{2}$, giving a maximum cost of $\max_{y\in X}q(y)(2c_{3}-2c_{2})$. However, the optimal facility placement is at $c_{3}$, which gives a maximum cost of $\min_{y\in X}q(y)(c_{3}-c_{2})$. Dividing these terms gives us the lower bound of $2r_{q}$.
To see that the approximation ratio is at most $2r_{q}$, note that a phantom mechanism with $p_{1}=0$ and $p_{n+1}=1$ ensures the facility is placed within $[x_{1},x_{n}]$, so the mechanism’s maximum cost is at most $(x_{n}-x_{1})\max_{y\in X}q(y)$, whilst the optimal maximum cost is at least $\frac{x_{n}-x_{1}}{2}\min_{y\in X}q(y)$. ∎
@@
Consider an agent location profile where $x_{1}=x_{2}=\dots=x_{n-1}$ and $x_{n}\neq x_{1}$. When the scaling function has its minimum value at $\frac{x_{1}+x_{n}}{2}$ and its maximum value on the facility/dictator, we have $\frac{MC(q,f(q,\boldsymbol{x}),\boldsymbol{x})}{MC(q,y^{*}_{MC},\boldsymbol{x})}=2r_{q}$, showing that the approximation ratio is at least $2r_{q}$. For the matching upper bound, we note that the mechanism’s maximum cost is at most $(x_{n}-x_{1})\max_{y\in X}q(y)$ and the optimal maximum cost is at least $\frac{x_{n}-x_{1}}{2}\min_{y\in X}q(y)$. ∎
@@ 8 Discussion
Acknowledgements
The authors would like to thank Matthias Greger for his extremely helpful comments and feedback. We also thank the Institute for Mathematical Sciences, National University of Singapore for facilitating our discussion with Matthias.